An association of City Mayors conducted a study to determine the average number of times

a family went to buy necessities in a week. They found that the mean is 4 times in a week. A

random sample of 20 families were asked and found a mean of 5 times in a week and a

standard deviation of 2. Use 5% significance level to test that the population mean is not

equal to 4. Assume that the population is normally distributed.


1
Expert's answer
2022-05-03T16:39:03-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=4"

"H_a:\mu\not=4"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=19" degrees of freedom, and the critical value for a two-tailed test is "t_c = 2.093024."

The rejection region for this two-tailed test is "R = \{t: |t| > 2.093024\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{X}-\mu}{s/\sqrt{n}}=\dfrac{5-4}{2/\sqrt{20}}\approx2.236068"

Since it is observed that "|t| = 2.236068 >2.093024=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=19" degrees of freedom, "t=2.236068" is "p=0.037541," and since "p=0.037541<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 4, at the "\alpha = 0.05" significance level.